OpenStudy (thats_kyy):

Lana makes $8 an hour plus $12 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. (3 points) Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. (3 points) Part C: Lana earned $368 in 1 week. How many hours (regular plus overtime) did she work? (4 points)

OpenStudy (s2rian):

part A ) W=8x

OpenStudy (s2rian):

part B ) S=8*40+12*(y-40)

OpenStudy (s2rian):

part C ) 368 is over 320 (4*80) so she works overtime 368=320+48 and 48 is 12*4. so she works 4 hours more than 40 hours

OpenStudy (thats_kyy):

I don't if it's correct is it?

OpenStudy (mathmale):

No one should be handing out answers without the involvement of the person who asked the questions. @s2rian, please take note. @Thats_Kyy : Please look at Part A: Letting wages be represented by W(x), W(x)=8x. Since the units of measurement of money here is in dollars. That $8 represents the rate of pay, which is $8/hour. Multiplying that by x (the number of hours worked) produces the total amount this person earned in x hours. What about Part B? Could you write an equation for wages, W(x) = ??? Note that someone who works overtime would be paid $8 per hour for the first 8 hours, and then (how much?) for the next y hours.

OpenStudy (thats_kyy):

Okay..

OpenStudy (mathmale):

If you'd prefer to calculate wages just by plugging in the total number of hours, then we need an expression that takes into account that the worker is paid $8/hour for the first 8 hours. Want more discussion, or are you satisfied?

OpenStudy (thats_kyy):

more disscusion

OpenStudy (thats_kyy):

i need too actually get it -.-

OpenStudy (mathmale):

OK. Now I'll listen to you as you explain what you still need to know.

OpenStudy (mathmale):

Regarding B: How much is this worker paid per hour for overtime work?

OpenStudy (thats_kyy):

8 dollas

OpenStudy (mathmale):

that's his regular (not overtime) rate of pay. The pay rate is higher for overtime work. Can you find this information in the problem statement?

OpenStudy (thats_kyy):

no

OpenStudy (mathmale):

Here are some examples: If the worker works 8 hours, she earns $8 per hour times 8 hours, or $64. I've just gone back to re-read the problem statement. Here's what it says: "Lana makes $8 an hour plus $12 an hour for every hour of overtime." So, if Lana works 8 hours, she earns $64. But if she works an additional hour (which means 1 hour of overtime), she earns the base pay of $64 plus $12 times 1 hour of overtime). What's the total?

OpenStudy (thats_kyy):

76.

OpenStudy (mathmale):

Good. Now suppose she works 8 hours plus two hours overtime. How much will she earn, total? Please show how you calculated your answer.

OpenStudy (thats_kyy):

64+12+12=88

OpenStudy (mathmale):

Really good. You've got it. Now, we need an equation for the wages of someone who works overtime whereby we input the total number of hours worked and the formula automatically subtracts 8 hours from that total time. We did that in the last 2 examples. Can you write such an equation? W(x) = $8*8 + ????

OpenStudy (mathmale):

Recall: If this woman worked 10 hours, 8 hours were at her standard pay of $8/hour and 2 hours were at her overtime rate of $12/hour. We want a formula W(x) that can handle this. We should be able to type in x = 10 and get the result $88. How would you write such an equation, where 8 hours is subtracted from the total number of hours worked?